OpenStudy (anonymous):

The numerator of a fraction is 7 times the denominator. If 3 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 60/12. Find the original fraction.

OpenStudy (anonymous):

d = denominator, n = numerator. n = 7d, and if n-3/d+3 = 60/12 then n/d = 63/9. also 9*7 = 63, so we know it's right.

OpenStudy (anonymous):

7x/x is first fraction 7x-3/x+3 after adding three fraction becomes 7x-3/x-3=60/12=5 7x-3=5x+15 2x=18 x=9 fraction is 63/9

OpenStudy (anonymous):

thank you :)